OpenStudy (anonymous):

A water ride at a local water park has a ride shaped like a cone that acts like a funnel whereby guests swirl around the cone until they drop through its center. There is one ride for adults and a similar, smaller version for children. If the adult ride has a radius of 25 feet and the child ride has a radius of 10 feet, what is the ratio between the volumes of each ride?

OpenStudy (callisto):

volume ratio of children : volume ratio of children = (10/25)^3 = 8/125 = 8:125

OpenStudy (anonymous):

perfect, I also got 125:8, I wanted to see if I was right